### v2

# Exploring the flow harmonic correlations via multi-particle symmetric and asymmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV

Abstract
We study multi-particle azimuthal correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. We use initial conditions obtained from a Monte-Carlo Glauber model and evolve them within a viscous relativistic hydrodynamics framework that eventually gives way to a transport model in the late hadronic stage of the evolution. We compute the multi-particle symmetric and asymmetric cumulants and present the results for their sensitivity to the shear and bulk viscosities during the hydrodynamic evolution. We also check their sensitivity to resonance decay and hadronic interactions. We show that these observables are more sensitive to the transport properties than the traditional flow observables.
preprint: APS/123-QED
I Introduction
In ultra-relativistic heavy-ion collisions, a thermalized, strongly interacting, deconfined medium, known as Quark-Gluon Plasma (QGP), is expected to form [1, 2, 3, 4, 5, 6]. The information regarding certain properties of QGP is carried by anisotropies observed in the momentum distributions of the produced particles [7, 8, 3]. After hadronization, particles are emitted anisotropically in the transverse plane to the beam direction, as detected by a detector. Traditionally, Fourier series in flow amplitudes $v_{n}$ and symmetry planes $\psi_{n}$ are used to quantify this azimuthal distribution in the momentum space [9],
$\frac{dN}{d\phi}=\frac{1}{2\pi}\left[1+2\sum_{n=1}^{\infty}v_{n}cos\left[n\left(\phi-\psi_{n}\right)\right]\right]$ (1)
Collective flow is crucial for studying the properties of the QGP medium formed in these collisions [10, 7, 11]. The impact parameter driven spatial geometry of the fireball, event-by-event nucleon position as well as sub-nucleonic fluctuations within the overlap region of the two colliding nuclei lead to the development of various orders of anisotropic flow harmonics. Due to the almond shape of the overlap region, the primary source of the second-order flow harmonic, known as elliptic flow ($v_{2}$), is the initial spatial anisotropy [12, 13, 14, 15, 16, 17, 18]. The third-order flow harmonic, known as triangular flow ($v_{3}$), arises from event-by-event fluctuations in the positions of the colliding nucleons as well as their sub-nucleonic fluctuations [19, 20, 21, 22, 23]. These fluctuations also give rise to higher-order flow harmonics. Previous studies have demonstrated that these flow harmonics are sensitive to the equation of state (EoS) and the transport properties of the fireball created in the collision, such as shear and bulk viscosity [24, 25, 26, 27, 28].
Anisotropic flow analysis involves measuring $v_{n}$, $\psi_{n}$, and their event-by-event correlations and fluctuations. Due to the random fluctuations of the impact parameter vector, however, the estimation of $v_{n}$ and $\psi_{n}$ involves the complications associated with reaction plane dispersion in conventional flow analyses. As a result, they are indirectly estimated using correlation techniques, in which there is no need for an event-by-event determination of the reaction plane [29, 30]. Therefore, there is no need to correct for dispersion in an estimated reaction plane. The foundation of this alternative approach is based on the following outcome,
$\langle e^{i\left(n_{1}\phi_{1}+n_{2}\phi_{2}+...+n_{k}\phi_{k}\right)}\rangle=v_{1}v_{2}...v_{k}e^{i\left(n_{1}\psi_{1}+n_{2}\psi_{2}+...+n_{k}\psi_{k}\right)}$ (2)
which analytically relates multiparticle azimuthal correlators and flow degrees of freedom [31]. The average is calculated over all unique sets of $k$ different particles in a single event. This expression can be used to determine the properties of flow amplitudes $v_{n}$ and symmetry planes $\psi_{n}$ on an event-by-event basis. Apart from collective flow, other sources of correlations called non-flow are also present that typically involve only a subset of particles.
Multivariate cumulants were introduced in anisotropic flow analyses in the early 2000s in Refs. [32, 33], which tackled long-standing issues in the field and transformed the approach to anisotropic flow analysis in high-energy physics. For example, the four-variate cumulant that is used to estimate the flow amplitude, $v_{n}$, from four-particle correlation is defined as [33]
$c_{n}\{4\}=\langle\langle e^{in(\phi_{1}+\phi_{2}-\phi_{3}-\phi_{4})}\rangle\rangle-2\langle\langle e^{in(\phi_{1}-\phi_{3})}\rangle\rangle^{2},$ (3)
The double angular brackets indicate that the averaging procedure is performed in two steps, averaging over all distinct particle multiplets in an event and then averaging these single-event averages with appropriately chosen event weights. By generalizing this concept for non-identical harmonics, new observables, which strictly satisfy all defining mathematical properties of cumulants, were constructed to quantify the correlations among different flow harmonics [34]. Ref. [35] further generalized these observables to probe the genuine correlation between different moments of flow harmonics.
I.1 Symmetric Cumulants of Flow Amplitudes
Reference [34] proposed a general algorithm to measure multi-particle correlation where the harmonics inside the correlator bracket of Eq. (3) can be different. This introduces a new set of observables, known as symmetric cumulants $(SCs)$, which can be used to measure the correlation between event-by-event fluctuation of flow harmonics $v_{m}$ and $v_{n}$. This new approach allows for the separation of non-flow and flow contributions and provides initial insights into how the combinatorial background contributes to flow measurements using correlation techniques. These four-particle symmetric cumulants, ${\rm SC}(m,n)$, are defined as [36]
${\rm SC}(m,n)$$\equiv$ $\cos(m\varphi_{1}\!+\!n\varphi_{2}\!-\!m\varphi_{3}-\!n\varphi_{4})\right>\right>_{c}$ (4)
$=$ $\cos(m\varphi_{1}\!+\!n\varphi_{2}\!-\!m\varphi_{3}-\!n\varphi_{4})\right>\right>$
[…]
The subscript $c$ indicates the cumulant. Positive (Negative) values of ${\rm SC}(m,n)$ suggest the correlation (anti-correlation) between $v_{m}^{2}$ and $v_{n}^{2}$, which means that if $v_{m}^{2}$ is larger than $\langle v_{m}^{2}\rangle$ in an event then the probability of $v_{n}^{2}$ being larger than $\langle v_{n}^{2}\rangle$ in that same event is enhanced (suppressed). The ${\rm SC}(m,n)$ observables focus on the correlations among different orders of flow harmonics and enable the quantitative comparison between experimental data and model calculations. To eliminate the effect of the magnitudes of $v_{m}$ and $v_{n}$ on the value of the symmetric cumulant, we divide ${\rm SC}(m,n)$ by their average values, $v_{m}^{2}\right>$ and $v_{n}^{2}\right>$, and define the normalized symmetric cumulant. This enables us to compare data and model calculations in a quantitative way and compare the fluctuations of the initial and final states. The normalized symmetric cumulant, denoted by ${\rm NSC}(m,n)$, is achieved following the standard method from Ref. [37]:
${\rm NSC}(m,n)=\frac{{\rm SC}(m,n)}{\langle v_{m}^{2}\rangle\langle v_{n}^{2}\rangle}.$ (5)
These correlations have been measured by both STAR experiment at RHIC [38] and ALICE experiment at LHC [39] with the observation of a positive correlation between $v_{2}$ and $v_{4}$ while a negative correlation between $v_{2}$ and $v_{3}$. The sensitivity of these observables to the shear viscosity to entropy density ($\eta/s$) has been studied in transport models like AMPT [40].
I.2 Asymmetric Cumulants of Flow Amplitudes
Recent studies indicate that insightful information can be obtained about the properties of Quark-Gluon Plasma by using higher-order observables [41]. These observables can probe the genuine correlations between the different moments of different flow harmonics. They are robust against non-flow correlations, which can be verified using the HIJING Monte Carlo generator [42]. The asymmetric cumulant ${\rm AC}_{2,1}(m,n)$ is defined as [35]
$\begin{split}{\rm AC}_{2,1}(m,n)&\quad\equiv\langle(v_{m}^{2})^{2}v_{n}^{2}\rangle_{c}\quad\equiv\langle v_{m}^{4}v_{n}^{2}\rangle_{c}\\ &\quad=\langle v_{m}^{4}v_{n}^{2}\rangle-\langle v_{m}^{4}\rangle\langle v_{n}^{2}\rangle-2\langle v_{m}^{2}v_{n}^{2}\rangle\langle v_{m}^{2}\rangle\\ &\quad+2\langle v_{m}^{2}\rangle^{2}\langle v_{n}^{2}\rangle\end{split}$ (6)
In Eq. (6), the subscript (2,1) on the left-hand side indicates the exponents of $v_{m}^{2}$ and $v_{n}^{2}$ on the right-hand side. The combinations of azimuthal correlators used to estimate the $ACs$ are [35]:
$\begin{split}{\rm AC}_{2,1}(m,n)&=\langle\langle{\rm e}^{i(m\varphi_{1}+m\varphi_{2}+n\varphi_{3}-m\varphi_{4}-m\varphi_{5}-n\varphi_{6})}\rangle\rangle\\ &\quad-\langle\langle{\rm e}^{i(m\varphi_{1}+m\varphi_{2}-m\varphi_{3}-m\varphi_{4})}\rangle\rangle\langle\langle{\rm e}^{i(n\varphi_{1}-n\varphi_{2})}\rangle\rangle\\ &\quad-2\langle\langle{\rm e}^{i(m\varphi_{1}+n\varphi_{2}-m\varphi_{3}-n\varphi_{4})}\rangle\rangle\langle\langle{\rm e}^{i(m\varphi_{1}-m\varphi_{2})}\rangle\rangle\\ &\quad+2\langle\langle{\rm e}^{i(m\varphi_{1}-m\varphi_{2})}\rangle\rangle^{2}\langle\langle{\rm e}^{i(n\varphi_{1}-n\varphi_{2})}\rangle\rangle\end{split}$ (7)
These expressions are genuine multivariate cumulants. ${\rm AC}_{1,1}(m,n)$ corresponds to SC$(m,n)$, illustrating the generalisation aspect of the $ACs$. We can also normalize the $ACs$. This procedure has two benefits. Normalizing the results allows proper comparisons and determination of the initial state effects and the changes brought by the hydrodynamic evolution since the predictions for $ACs$ in the initial and final state do not have the same scale. Flow amplitudes depend on the transverse momentum, $p_{T}$, which leads to a similar dependence in any linear combinations of them, such as the $SCs$ and the $ACs$. The normalization eliminates this dependence and enables comparisons between models and data with different $p_{T}$ ranges [42]. Again, the normalization of the $ACs$ is done following the standard method from Ref. [37],
${\rm NAC}_{2,1}(m,n)$ $=$ $\frac{{{\rm AC}}_{2,1}(m,n)}{\langle v_{m}^{2}\rangle^{2}\langle v_{n}^{2}\rangle}.$ (8)
In this paper, we have presented the measurement of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$ in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV using a hybrid hydrodynamic model. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 42, 37, 44, 45, 46, 47], while ${\rm AC}_{2,1}(m,n)$ have been computed in transport models [45, 46]. ALICE collaboration has measured ${\rm AC}_{2,1}(m,n)$ for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [48]. In addition to symmetric cumulants, various other correlators have also been studied [49, 50, 51, 52]. We show the sensitivity of these measurements to the transport properties, such as the shear viscosity to entropy density ratio, $\eta/s$, and the bulk viscosity to entropy density ratio, $\zeta/s$, of the medium.
II Framework
We use a framework with multiple components to simulate various stages of heavy ion collisions. The hydrodynamic evolution has been initialized using a Glauber-based model. To generate the incoming nuclei, we use a Woods-Saxon distribution to sample nucleons.
$\rho=\frac{\rho_{0}}{1+exp[(r-R)/a]}.$ (9)
[…]
The evolution of the energy-momentum tensor starts at $\tau_{0}=0.6$ fm. The hydrodynamic equations are solved using the Music simulation [54, 55, 56, 57], which utilizes a Kurganov-Tadmor algorithm. A constant effective $\eta/s$ = 0.08 was fixed by reproducing the measured anisotropic flow coefficients of charged hadrons. We use a temperature-dependent specific bulk viscosity parametrized in the following way [58]:
$\frac{\zeta}{s}=\left\{\begin{array}[]{ll}\lambda_{1}e^{-(x-1)/\sigma_{1}}+\lambda_{2}e^{-(x-1)/\sigma_{2}}+0.001&(T>1.05T_{C})\\ \lambda_{3}e^{(x-1)/\sigma_{3}}+\lambda_{4}e^{(x-1)/\sigma_{4}}+0.03&(T<0.995T_{C})\end{array}\right.$ (10)
where $x=T/T_{C}$. The fitted parameters are $\lambda_{1}=\lambda_{3}=0.9$ , $\lambda_{2}=0.25$, $\lambda_{4}=0.22$, $\sigma_{1}=0.025$, $\sigma_{2}=0.13,$ $\sigma_{3}=0.0025$, $\sigma_{4}=0.022,$ $A_{1}=-13.77$, $A_{2}=27.55$ and $A_{3}=13.45$.
We use a QCD equation of state (EoS), NEoS-B [59], based on continuum extrapolated lattice calculations at zero net baryon chemical potential published by the HotQCD Collaboration [60, 61, 62]. It is smoothly matched to a hadron resonance gas EoS in the temperature region between 110 and 130 MeV [63].
A hyper-surface is generated from the hydrodynamic space-time evolution of the fluid. To describe the dilute hadronic phase, we utilize the iSS code [64, 65] to sample the primordial hadrons from the hypersurface with a constant energy density, $e_{sw}=0.26\ GeV/fm^{3}$, equivalent to a local temperature of approximately 151 MeV. Then, we use the UrQMD code [66, 67] to simulate scatterings and decays of hadrons during the late stage of heavy ion collisions. To increase statistics, each hydrodynamic switching hyper-surface is sampled multiple times. The number of oversampling events for every hydrodynamic event is determined to ensure sufficient statistics for every centrality class. We generated ensembles of 2000 hydrodynamic events for each centrality class, each with a suitable number of repeated samplings. Table 1 lists the number of events generated for each of the centrality classes with the three sets of parameters for the hydrodynamic stage shown in Table 2. Table 3 outlines the particles used in the analyses at the end of different stages of the evolution. Analysis at the end of:
$\bullet$ stage I includes hydrodynamic evolution and particlization,
$\bullet$ stage II includes hydrodynamic evolution, particlization, and resonance decay, and
$\bullet$ stage III includes hydrodynamic evolution, particlization, resonance decay, and hadronic interactions within the UrQMD model approach.
Centrality class 0-10% 10-20% 20-30% 30-40% 40-50%
[…]
Table 1: Ensemble size (in millions) of the centrality classes at the end of Stage III.
$\eta/s$ $\zeta/s$
Parameter(Par.) Set I 0 0
Parameter(Par.) Set II 0.08 0
Parameter(Par.) Set III 0.08 $\zeta/s(T)$
Table 2: The transport coefficients’ values for the parameter sets used in the ensembles generated for this study.
Particle list
Stage I Hydrodynamic evolution + particlization (Primordial)
Stage II Primordial + Resonace decays
Stage III Primordial + UrQMD
Table 3: The particle lists analyzed at the end of different stages of the evolution (Hydrodynamic evolution, resonance decay, and hadronic transport).
In the following results, some observables will be presented by scaling with the average number of participants, $\langle N_{part}\rangle$. Table 4 lists the values of $\langle N_{part}\rangle$ in different centrality bins, which are determined using the Monte-Carlo Glauber model. The ${\rm SC}(m,n)$ and ${\rm AC}(m,n)$ observables have been computed using particles with transverse momentum ($p_{T}$) ranges of [0.2, 2.0] GeV and [0.01, 3.0] GeV, respectively.
Centrality class $\langle N_{part}\rangle$
0-10% 327.7
10-20% 240.3
20-30% 172.7
30-40% 120.5
40-50% 81.5
Table 4: $\langle N_{part}\rangle$ in the centrality classes from a Monte-Carlo Glauber model.
III Results
We calibrated our model by using the produced particle yields, transverse momentum spectra and elliptic flow. We then compared these results to experimental data from the PHENIX and STAR Collaborations. The normalization factor for the system’s energy density was determined by matching the model and STAR data [68] charged hadron multiplicity $dN_{Ch}/d\eta$ in the 0-5% centrality bin. We also compared the $p_{T}$ spectra of identified particles in 0-5% centrality Au+Au collisions to STAR results [68]. Additionally, we compared the single-particle anisotropic flow coefficient $v_{2}$ of charged particles in 20-30% centrality Au+Au collisions with the PHENIX measurement [69]. The hydrodynamic evolution of the MCGM initial conditions, coupled to UrQMD, can reproduce the elliptic flow up to the mid-central collisions.
III.1 Sensitivity to transport coefficients
III.1.1 Four-particle symmetric cumulants
In this section, we present results for ${\rm SC}(m,n)$ from our simulations. To check their sensitivity to the hydro transport coefficients, we compute these observables for the three sets of ensembles mentioned in Table  2. We observe that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by shear and bulk viscosities. In Fig.[1], we show the effect of shear viscosity on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. In $v_{2}\{4\}$, we see a maximum change of around 10% in the mid-central collisions. While in ${\rm SC}(2,3)$, we see a change of around 40-50% throughout the centrality range. We observe a similar dependence on the bulk viscosity of the medium. Thus, these observables are more sensitive to the transport coefficients. Similar to ${\rm SC}(2,3)$, ${\rm SC}(2,4)$ is also affected by the transport coefficients in the same manner. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals a similar difference in their sensitivity to shear and bulk viscosity.
Figure 1: Ratios of $v_{2}\{4\}$ and ${\rm SC}(2,3)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm SC}(2,3)$ using lines and markers. The red line and the cyan band show the effects of $\eta/s$, and the blue line and the magenta band show the effects of $\zeta/s$.
III.1.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig.[2], we show the effect of shear viscosity on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,4)$. For ${\rm AC}_{2,1}(2,4)$, we see a change of around 50% throughout the centrality range (note that 0-10% and 10-20% centrality classes have large statistical uncertainties). We observe that the medium’s bulk viscosity also has a similar effect on these observables. Compared to the traditional flow observables, we see that these are more sensitive to the transport coefficients.
Figure 2: Ratios of $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$, with different hydro parameter sets, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of $\eta/s$ and $\zeta/s$ on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm AC}_{2,1}(2,4)$ using lines and markers. The red line and the cyan band show the effects of $\eta/s$, and the blue line and the magenta band show the effects of $\zeta/s$.
III.2 Sensitivity to hadronic interactions
III.2.1 Four-particle symmetric cumulants
The hadronic afterburner plays an essential role in describing the data collected at RHIC [43]. To check their sensitivity to hadronic interactions and resonance decay, we compute these observables at the end of the three stages of the evolution mentioned in Table  3. We observe that both ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ are suppressed by resonance decay and enhanced by hadronic interactions. However, the magnitudes of these effects are different for the two observables. In Fig.[3], we show the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. In $v_{2}\{4\}$, we see a change of around 15% due to resonance decay and 5% due to hadronic interactions plus resonance decay in the mid-central collisions. While in ${\rm SC}(2,3)$, we see a change of around 40% due to resonance decay throughout the centrality range and 40-60% due to hadronic interactions plus resonance decay. ${\rm SC}(2,4)$ is also affected in the same manner. Thus, these observables are more sensitive to resonance decay and hadronic interactions. The comparison of the symmetric cumulants with $v_{3}\{4\}$ also reveals similar differences in their sensitivity to these effects.
Figure 3: Ratios of $v_{2}\{4\}$ and ${\rm SC}(2,3)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm SC}(2,3)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm SC}(2,3)$ using lines and markers. The red line and the cyan band show the effects of resonance decay, and the blue line and the magenta band show the effects of hadronic interactions.
III.2.2 Six-particle asymmetric cumulants
Now, we present the results for ${\rm AC}_{2,1}(m,n)$. In Fig.[4], we show the effects of resonance decay and hadronic interactions on elliptic flow from the four-particle correlations and ${\rm AC}_{2,1}(2,4)$. For ${\rm AC}_{2,1}(2,4)$, we see a change of around 40% due to these in the mid-central collisions. Although the mean values are somewhat similar throughout centrality, due to large uncertainties (about 2X), AC results for 0-10% and 10-20% centralities have been excluded for clarity. ${\rm AC}_{2,1}(2,3)$ is also affected in the same manner. Compared to the traditional flow observables, we see that these are more sensitive to resonance decay and hadronic interactions.
Figure 4: Ratios of $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$, computed at the three different stages of the evolution, vs centrality in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. The figure compares the effects of resonance decay and hadronic interactions on $v_{2}\{4\}$ and ${\rm AC}_{2,1}(2,4)$. The plotted results show ratios of $v_{2}\{4\}$ using bands and ${\rm AC}_{2,1}(2,4)$ using lines and markers. The red line and the cyan band show the effects of resonance decay, and the blue line and the magenta band show the effects of hadronic interactions. AC results for 0-10% and 10-20% centralities, with large uncertainties, have been excluded for clarity.
III.3 Centrality dependence of ${\rm SC}(m,n)$ and ${\rm AC}_{2,1}(m,n)$
III.3.1 Four-particle symmetric cumulants
Next, we compare the results from our simulations with the experimental measurements from RHIC. ${\rm SC}(m,n)$ have been computed in previous studies from hydrodynamic and transport models [43, 42, 37, 44, 45, 46, 47], while ${\rm AC}_{2,1}(m,n)$ have been computed in some transport models [45, 46]. Fig. [5] shows comparisons between the theory calculations for the symmetric cumulants ${\rm SC}(2,3)$ and ${\rm SC}(2,4)$ multiplied by $\langle N_{part}\rangle$ and the same measured by the STAR Collaboration [70]. The factor of $\langle N_{part}\rangle$ is multiplied to scale out the trivial dilution of correlation with the increase in the number of multiplets. We are able to reproduce the qualitative variations with centrality. The magnitudes of $\langle N_{part}\rangle{\rm SC}(2,3)$ and $\langle N_{part}\rangle{\rm SC}(2,4)$ increase from central to mid-peripheral events and then again decreases for very peripheral events. Here, we do not aim to reproduce the experimental data because our aim is to check the sensitivity of these observables to the transport properties of the system created in these collisions. The comparison in Fig.[5] reflects that the correlation in initial spatial eccentricities and the following hydrodynamic evolution can capture the correlation among flow harmonics of different orders. Negative values of ${\rm SC}(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as ${\rm SC}(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. Symmetric cumulants do not have contributions from non-flow effects, where non-flow refers to azimuthal correlations not related to the reaction plane orientation, like those from resonances, jets, quantum statistics, etc. This is verified by computing these observables for the HIJING model, which includes only non-flow physics, for which these are consistently zero [35]. The model captures the variation of ${\rm SC}(m,n)$ with the centrality. In central collisions, the model underestimates ${\rm SC}(2,4)$, possibly due to inadequate fluctuations in the initial conditions or some shortcomings in the transport model since the nonlinear response in $v_{4}$ and $\varepsilon_{4}$ is sensitive to these effects. However, ${\rm SC}(2,3)$ is better aligned with the data because the $v_{2}-v_{3}$ correlations mainly arise from the initial state $\varepsilon_{2}-\varepsilon_{3}$ correlations due to linear response in $v_{2}$ and $v_{3}$.
Figure 5: Symmetric cumulant vs centrality from four-particle correlations compared with STAR data [70] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm SC}(2,3)$ is consistently negative while as ${\rm SC}(2,4)$ is consistently positive.
The normalized symmetric cumulants in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV, computed using Eq.[5] in momentum space and the corresponding equation in coordinate space for the initial state eccentricity (Eq.[11]) evaluated using the Monte Carlo Glauber model, are presented in Fig.[6].
[…]
If only eccentricity drives $v_{n}$, then we can expect that the ${\rm NSC}(m,n)$ in the final state would be equal to the ${\rm NSC}(m,n)$ in the initial state. Fig.[6] shows that the initial anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ is mainly responsible for the observed anti-correlation between $v_{2}$ and $v_{3}$. However, the correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is smaller than the observed correlation between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ not only comes from the linear response of the system to $\varepsilon_{4}$ but also has a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NSC}(2,4)$ in the final state and the initial state increases, likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$. This has also been observed in Pb+Pb collisions at $\sqrt{s_{NN}}$ = 2.76 TeV by the ATLAS [71] and ALICE [39] experiments. The properties of the medium were suggested to affect the relative contribution of $\varepsilon_{2}$ in $v_{4}$ compared to $\varepsilon_{4}$ [72]. Therefore, ${\rm NSC}(2,4)$ provides a probe into the medium properties.
Figure 6: Normalized symmetric cumulant vs centrality from four-particle correlations compared with STAR data [70] in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm NSC}(2,3)$ is consistently negative while as ${\rm NSC}(2,4)$ is consistently positive.
III.3.2 Six-particle asymmetric cumulants
${\rm AC}_{2,1}(m,n)$ have measured by the ALICE collaboration for Pb+Pb collisions at $\sqrt{s_{NN}}=5.02\ TeV$ [48]. In Fig.[7], we should the centrality dependence of ${\rm AC}_{2,1}(m,n)$, multiplied by $\langle N_{part}\rangle$ to scale out the trivial dilution of correlation with the increase in the number of multiplets, for Au+Au at $\sqrt{s_{NN}}=200\ GeV$. The magnitudes of $\langle N_{part}\rangle{\rm AC}_{2,1}(2,3)$ and $\langle N_{part}\rangle{\rm AC}_{2,1}(2,4)$ increase with centrality. The trends are similar to the results by the ALICE collaboration. As for symmetric cumulants, negative values of $AC(2,3)$ throughout the centrality range reveal the anti-correlation between $v_{2}$ and $v_{3}$. While as $AC(2,4)$ is positive for all centralities, indicating positive correlations between $v_{2}$ and $v_{4}$. We compute these observables for the three sets of ensembles mentioned in Table  2 and observe that both ${\rm AC}_{2,1}(2,3)$ and ${\rm AC}_{2,1}(2,4)$ are suppressed by shear and bulk viscosities.
Figure 7: Asymmetric cumulant vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm AC}_{2,1}(2,3)$ is consistently negative while as ${\rm AC}_{2,1}(2,4)$ is consistently positive.
In Fig.[8], we have compared the normalized asymmetric cumulant in coordinate and momentum space, as computed using Eq.[8] and the corresponding equation for the initial state eccentricities (Eq.[11]). Again, if only eccentricity drives $v_{n}$, we can expect the ${\rm NAC}(m,n)$ in the final state to be equal to the ${\rm NAC}(m,n)$ in the initial state. Fig. [8] illustrates that the initial negative correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$ primarily causes the observed negative correlation between $v_{2}$ and $v_{3}$. However, unlike symmetric cumulants, here, the initial state correlations are smaller in the peripheral collisions. The correlation between $\varepsilon_{2}$ and $\varepsilon_{4}$ is weaker than the correlation observed between $v_{2}$ and $v_{4}$. The contribution to anisotropic flow $v_{4}$ comes not only from the linear response of the system to $\varepsilon_{4}$, but also from a contribution proportional to $\varepsilon_{2}^{2}$. As collisions become more peripheral, the difference between ${\rm NAC}(2,4)$ in the final state and the initial state increases. This is likely due to $\varepsilon_{2}$ playing a more significant role in $v_{4}$.
Figure 8: Normalized asymmetric cumulants vs centrality from six-particle correlations in Au+Au collisions at $\sqrt{s_{NN}}$ = 200 GeV. ${\rm NAC}_{2,1}(2,3)$ is consistently negative while as ${\rm NAC}_{2,1}(2,4)$ is consistently positive.
III.4 Conclusions
We have presented results for multi-particle correlation functions in heavy-ion collisions at top RHIC energy using a hybrid framework based on the Monte-Carlo Glauber model, MUSIC viscous hydrodynamics simulations, and the UrQMD hadronic cascade.
First, we adjusted the free parameters, such as shear and bulk viscosities, to describe particle multiplicities, mean transverse momentum, and anisotropic flow. After that, we discussed the impact of shear and bulk viscosities, resonance decay, and hadronic interactions on traditional and novel observables.
We have studied correlators that measure the correlations between flow harmonics of varying orders, specifically four and six-particle correlations, including both symmetric and asymmetric cumulants as functions of centrality at midrapidity in Au+Au collisions at $\sqrt{s_{NN}}=200$ GeV. These observables are very sensitive to shear and bulk viscosities, resonance decay, and hadronic interactions. A single parameter set of transport coefficients is unable to explain $SC(2,3)$ and $SC(2,4)$ simultaneously.
We observed anti-correlation between event-by-event fluctuations of $v_{2}$ and $v_{3}$, while the event-by-event fluctuations of $v_{2}$ and $v_{4}$ are found to be positively correlated. The observed anti-correlation between $v_{2}$ and $v_{3}$ appears to be described by the initial-stage anti-correlation between $\varepsilon_{2}$ and $\varepsilon_{3}$, which supports the idea of linearity between $\varepsilon_{n}$ and $v_{n}$ [23]. The hydrodynamic response suggests that the final state fluctuation originates from the initial state. However, including the nonlinear hydrodynamic response of the medium is necessary to reproduce the measured correlation between $v_{2}$ and $v_{4}$, as the initial-stage linear correlation alone is insufficient.
III.5 Acknowledgements
We would like to acknowledge Tribhuban Parida for fruitful discussions and providing us the computational setup.